Solve the

following word problems in kinematics. Show your complete solution!

1.     

A plane flies 955 km due east, then

turns to north and flies another 649 km. Determine the polar coordinates of the

plane’s final position.


2.     

The slowest animal ever discovered is a

crab found in the Red Sea that travels an average speed of 5.7 km/yr. How many

days will it take this crab to travel 1 meter?


3.     

A feather is dropped on the moon from a

height of 1.80 meters. The acceleration of gravity on the moon is 1.67 m/s^2.

Determine the time for the feather to fall to the surface of the moon.

4.     

An airplane accelerates down a runway

from rest at 2.65 m/s^2 for 36.86 s until is finally lifts off the

ground. Determine the distance traveled before takeoff.




1
Expert's answer
2021-10-14T11:36:53-0400

1) "\rho=\sqrt{955^2+649^2}\approx1155\ (km)"


"\alpha=\tan^{-1}(649/955)=34.2°"


2) "t=s/v=1/(5700/365)=0.064\ (days)"


3) "h=gt^2/2\to t=\sqrt{2h/g}=\sqrt{2\cdot1.8/1.67}=1.47\ (s)"


4) "s=at^2/2=2.65\cdot36.86^2/2=1800 \ (m)"